what is the difference between
6^2 * 6^2/6 and 6^2 + 6^2/6 ?

Answer by rothauserc(4718)

6^2 * 6^2/6 = 6^(12/6) * 6^(2/6) = 6^(14/6)
we add exponents of like base when they are multiplied
6^(14/6) - 6^(12/6) - 6^(2/6)
reduce fractional exponents
6^(7/3) - 6^2 - 6^(1/3)
note that 6^(7/3) = 3 root of 6^7 and 6^(12/6) = 6^2
65.416341342 - 36 - 1.817120593 = 27.599220749